3x + 5y -z =11
2x + y - 3z=12
y + 2z = -4

Answer by mananth(16946)

3x+5y-z =11 -------------- 1
2x+y-3z =12 -------------- 2
y+2z=-4 -------------- 3

consider equation 1 &2 Eliminate y
Multiply 1 by 1
Multiply 2 by -4
we get
3x+5y-z = 11
-8x+y+12z= -48
Add the two
-5x+6y+11z= -37 ------------- 4
consider equation 2 & 3 Eliminate y
Multiply 2 by -1
Multiply 3 by 1
we get
-2 x + -1 y + 3 z = -12
0 x + 1 y + 2 z = -4
Add the two
-2x+5z= -16 -------------5
Consider (4) & (5)Eliminate x
Multiply 4 by 2
Multiply (5) by -5
we get 2
-10 x + 22 z = -74
10 x + -25 z = 80
Add the two
-3 z = 6
/ -3
z = -2
Plug the value of z in 5
-2 x + -10 = -16
-2 x = -6
x = 3
plug value of x & z in 1
9 + 5 y + 2 = 11
5 y = -9 + -2 + 11
y= 0

m.ananth@hotmail.ca